OpenStudy (anonymous):

The function for the cost of materials to make a shirt is f(x) = five sixths x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts. 106 120 185 196

OpenStudy (anonymous):

i think its 196 but im not sure

OpenStudy (nfcfox):

Well they give us 18 shirts and we know that x represents amount of shirts and we want to plug that value into g(x) So we plug 18 into the f(x)

OpenStudy (nfcfox):

5/6(18)+5= 15+5=20 Now we plug 20 into g(x). 5(20)+6= 100+6= 106. The answer is 106